# sample.bib

### v2

@@ 4.3 Demand Drivers and Opportunities Identification
Through the comprehensive activities conducted under the collaboration components outlined in Section 4.1, the core team engaged in extensive analysis with network participants, complemented by a detailed evaluation of Polish national documents and related reports. This effort aimed to elucidate the diverse demand drivers for CCS/CCUS deployment in Poland, while simultaneously capturing the perspectives of Norwegian industrial and R&D stakeholders.
A primary driver identified is the civil society’s expectation for a transition towards a future characterized by net-zero emissions. Government policy and regulations are increasingly reflecting this societal demand, thereby strengthening the business case for investing in CCS technologies. CCS is recognized as a vital instrument for reducing CO2 emission exposure in specific industrial sectors, presenting a dual opportunity to minimize greenhouse gases while fostering a burgeoning industry. It is positioned within scenarios as one of the most efficient pathways to achieve net-zero targets, ensuring a just transition for industrial and highly carbon-dependent communities.
@@
Figure 4: Interplay of governing factors and opportunities influencing the CCS deployment in Poland. Nodes represent key elements required to achieve emission reductions goals. Directed edges illustrate the influences and dependencies among these factors, forming a comprehensive framework to navigate complexities in adaptation and governance. This visualization highlights critical leverage points, policy implications, and strategies for effective stakeholder engagement, serving as a pivotal tool for constructing a roadmap for bilateral Norwegian-Polish cooperation and advancing CCS deployment efforts in Poland.
@@
The regulatory environment remains in a nascent state, with numerous aspects yet to provide assured pathways for CCS deployment. While funding support is accessible through European Union channels, including Programs of Common Interest favoring large-scale regional infrastructural endeavors, stakeholder hesitations persist. These apprehensions stem from the high costs, unclear government support, financing difficulties, and complex administrative protocols involved in implementation. A broader challenge is the lack of public and institutional awareness regarding CCS technologies, which the Advanced Demand Drivers and Opportunities diagram seeks to address through its emphasis on strategic public outreach and policy integration frameworks.
@@ 5.1 Awareness, attitudes, and perception
The surveys highlight that a significant portion of Polish respondents are aware of the detrimental effects of carbon dioxide on atmospheric conditions. A substantial 78% acknowledge the critical role of CO2 emissions in climate change, with 77% supporting the reduction of emissions, including CO2 and other greenhouse gases, to preserve environmental health. Climate change poses a considerable challenge, recognized by 75% of respondents, and there is a perceived need for investment in CO2 capture technologies, with 73% expressing this viewpoint. Despite recognizing governmental endeavors in emission reduction, 57% find these efforts inadequate and advocate for new solutions, although only 40% show awareness of CCS methods as a viable remedy.
@@ 5.2 Implementation challenges and outlook
CCS garners support for installation near industrial zones, but its application near residential areas reveals mixed sentiments. Industrial CO2 capture facilities receive a favorable reception, yet underground storage invites varied responses, with 28% approval and 32% opposition. Concerns are fuelled by insufficient information and fears of unspecified risks, including possible land expropriation (52%), increased heavy transport traffic (46%), and CO2 explosion threats (40%), underlining educational gaps. Local impacts such as noise (33%) and visual landscape changes (31%) are seen as less problematic.
@@
Furthermore, Polish-Norwegian cooperation in CCS is favorably regarded, with 58% perceiving more advantages than drawbacks, a rise from 53%, and reduced indecision to 20% from 26%. Awareness of mutual benefits, such as Norway receiving Poland’s CO2, increased to 39% from 35%. The reduced perception of asymmetric benefits for Poland and the increased belief in the balanced advantages of Norwegian decarbonization involvement highlights a promising collaborative landscape.
@@ 6.1 R&D and Academic Training
Joint Research and Development (R&D) activities have emerged as critical components for the success and sustainability of the Norwegian-Polish CCS partnership. These collaborations focus strategically on site-specific considerations essential for facilitating CO2 storage pilots and full-scale sequestration within Poland’s geological formations. Innovation in carbon capture techniques, the advancement of existing technologies, and cooperative testing or piloting at facilities like the Technology Center Mongstad in Norway are proposed to align with local industry requirements.
@@
The academic and R&D sectors are integral to any pilot or industrial studies, particularly where technological or field-related questions arise [nooraiepour2018rock]. These questions necessitate collaboration with competent researchers to delve into and support industrial endeavors. Moreover, these sectors can significantly contribute to studies related to hub formation, technology retrofitting, geological-site subsurface evaluations, transport planning, carbon accounting, and more, thereby enhancing the overall efficacy and breadth of CCS initiatives.
@@ 6.2 Industrial synergies and joint ventures
Through this collaboration, opportunities for industrial synergies and technology transfer between Norwegian and Polish entities abound. Sharing expertise in project operations, maintenance, and joint development enhances Polish industries’ capabilities, enabling them to leverage Norwegian insights into CCS technology while mitigating associated risks. Establishing joint ventures and partnerships accelerates project deployment and opens new business avenues, aligning both countries with rising CCS facility demand.
A cornerstone of the initiative is the inclusive and well-rounded development of infrastructure necessary for the effective transport and storage of CO2 across Poland. Conducting comprehensive feasibility assessments of various CO2 transport modalities—such as pipelines, railways, inland waterways, and export ports for shipping to international facilities like the Northern Lights project in the North Sea—is paramount. The potential for repurposing existing pipelines or situating new routes along current corridors offers opportunities for optimizing infrastructure usage. Detailed feasibility studies, carried out in collaboration with the research and development (R&D) sector, assess the technical viability and explore a range of scenarios. This integration of diverse solutions forms a key component of the national CCS strategy, ensuring that infrastructure development is both strategic and forward-looking to support Poland’s climate mitigation goals.
@@ 6.3 Public engagement and stakeholder participation
Active public engagement and participation, identified in stakeholder workshops, are critical elements for compelling CCS deployment. Education, awareness, and fighting misinformation are vital for addressing concerns and building trust. Facilitating open dialogues with local communities and stakeholders fosters participation, resolves potential issues, and solidifies public support, a cornerstone of sustainable implementation. Drawing from Norway’s experiences, Poland can adopt effective public communication practices to boost confidence in CCS solutions.
@@ Acknowledgments
The authors acknowledge the support and funding received from the "Polish-Norwegian CCS Network: Acceleration of Climate Change Mitigation Technologies Deployment" project, funded through the EEA and Norway Grants under the Green Transition Collaborations Programme. This research was facilitated by the collaborative efforts between the AGH University of Krakow, University of Oslo, and Norwegian Energy Partners (NORWEP).

### v3

@@ 4.3 Demand Drivers and Opportunities Identification
Through the comprehensive activities conducted under the collaboration components outlined in Section 4.1, the core team engaged in extensive analysis with network participants, complemented by a detailed evaluation of Polish national documents and related reports. This effort aimed to elucidate the diverse demand drivers for CCS/CCUS deployment in Poland while simultaneously capturing the perspectives of Norwegian industrial and R&D stakeholders.
A primary driver identified is civil society’s expectation of a transition towards a future characterized by net-zero emissions. Government policy and regulations are increasingly reflecting this societal demand, thereby strengthening the business case for investing in CCS technologies. CCS is recognized as a vital instrument for reducing CO2 emission exposure in specific industrial sectors, presenting a dual opportunity to minimize greenhouse gases while fostering a burgeoning industry. It is positioned within scenarios as one of the most efficient pathways to achieve net-zero targets, ensuring a just transition for industrial and highly carbon-dependent communities.
@@
Figure 5: Interplay of governing factors and opportunities influencing the CCS deployment in Poland. Nodes represent key elements required to achieve emission reduction goals. Directed edges illustrate the influences and dependencies among these factors, forming a comprehensive framework to navigate complexities in adaptation and governance. This visualization highlights critical leverage points, policy implications, and strategies for effective stakeholder engagement, serving as a pivotal tool for constructing a roadmap for bilateral Norwegian-Polish cooperation and advancing CCS deployment efforts in Poland.
@@
The regulatory environment remains nascent, with numerous aspects yet to provide assured pathways for CCS deployment. While funding support is accessible through European Union channels, including Programs of Common Interest favoring large-scale regional infrastructural endeavors, stakeholder hesitations persist. These apprehensions stem from the high costs, unclear government support, financing difficulties, and complex administrative protocols involved in implementation. A broader challenge is the lack of public and institutional awareness regarding CCS technologies, which the Advanced Demand Drivers and Opportunities diagram seeks to address through its emphasis on strategic public outreach and policy integration frameworks.
@@ 5 Public Opinion towards CCS in Poland
Figure 6: Distribution and insights from pre- and post-campaign surveys, displaying multiple facets of Polish public perception and attitudes regarding environmental awareness and CCS technology. Subplots show: (a) public awareness levels about CO2 emissions and technology investments required, (b) sentiments and safety perceptions about CCS before and after an informational campaign, (c) engagement with various information sources, (d) concerns regarding the practical implementation of CCS projects, (e) perceptions and views on international cooperation related to CCS, particularly with Poland and Norway, and (f-h) changes in awareness and sentiment as well as source engagement before and after information campaigns. The results highlight significant perceptual shifts post-campaign and indicate major areas of public concern and interest.
@@ 5.1 Awareness, attitudes, and perception
The surveys highlight that a significant portion of Polish respondents are aware of the detrimental effects of carbon dioxide on atmospheric conditions. As Figure 6 demonstrates, a substantial 78% acknowledge the critical role of CO2 emissions in climate change, with 77% supporting the reduction of emissions, including CO2 and other greenhouse gases, to preserve environmental health. Climate change poses a considerable challenge, recognized by 75% of respondents, and there is a perceived need for investment in CO2 capture technologies, with 73% expressing this viewpoint. Despite recognizing governmental endeavors in emission reduction, 57% find these efforts inadequate and advocate for new solutions, although only 40% show awareness of CCS methods as a viable remedy (Fig. 6).
@@ 5.2 Implementation challenges and outlook
CCS garners support for installation near industrial zones, but its application near residential areas reveals mixed sentiments. Industrial CO2 capture facilities receive a favorable reception, yet underground storage invites varied responses, with 28% approval and 32% opposition. Concerns are fuelled by insufficient information and fears of unspecified risks, including possible land expropriation (52%), increased heavy transport traffic (46%), and CO2 explosion threats (40%), underlining educational gaps. Local impacts such as noise (33%) and visual landscape changes (31%) are seen as less problematic (Fig. 6).
@@
Furthermore, as shown in Figure 6, Polish-Norwegian cooperation in CCS is favorably regarded, with 58% perceiving more advantages than drawbacks, a rise from 53%, and reduced indecision to 20% from 26%. Awareness of mutual benefits, such as Norway receiving Poland’s CO2, increased to 39% from 35%. The reduced perception of asymmetric benefits for Poland and the increased belief in the balanced advantages of Norwegian decarbonization involvement highlights a promising collaborative landscape.
@@ 6.1 R&D and Academic Training
Joint Research and Development (R&D) activities have emerged as critical components for the success and sustainability of the Norwegian-Polish CCS partnership. These collaborations focus strategically on site-specific considerations essential for facilitating CO2 storage pilots and full-scale sequestration within Poland’s geological formations. Innovation in carbon capture techniques, the advancement of existing technologies, and cooperative testing or piloting at facilities like the Technology Center Mongstad (refer to (Fig. 3) in Norway are proposed to align with local industry requirements.
@@
The academic and R&D sectors are integral to any pilot or industrial studies, particularly where technological or field-related questions arise. These questions necessitate collaboration with competent researchers to delve into and support industrial endeavors. Moreover, these sectors can significantly contribute to studies related to hub formation, technology retrofitting, geological-site subsurface evaluations, transport planning, carbon accounting, and more, thereby enhancing the overall efficacy and breadth of CCS initiatives.
@@ 6.2 Industrial Synergies and Joint Ventures
Through this collaboration, opportunities for industrial synergies and technology transfer between Norwegian and Polish entities abound. Sharing expertise in project operations, maintenance, and joint development enhances Polish industries’ capabilities, enabling them to leverage Norwegian insights into CCS technology while mitigating associated risks. Establishing joint ventures and partnerships accelerates project deployment and opens new business avenues, aligning both countries with the rising demand for CCS facilities.
Besides joint industrial collaboration for developing capture hubs and providing fit-for-purpose capture technology services and consultations to emitters, an essential aspect is facilitating swift internalization of emission reduction requirements. CO2 storage collaborations emerge as a critical focal area. Until joint efforts, along with contributions from the R&D sector as mentioned in an earlier section, yield results through necessary pre-assessments, site-specific characterization, and pilots in Poland saline aquifers and hydrocarbon fields. The available storage pore volume in the Norwegian Continental Shelf (NCS), under recently awarded North Sea storage licenses (Fig. 3), can serve as an immediate and viable option for transporting and storing Polish carbon emissions.
Moreover, developments in the Gdansk area aimed at establishing a terminal (Fig.2) can align seamlessly with the Long Ship and similar initiatives to jumpstart CCS deployment. This coordination ensures that captured CO2 can be safely and readily stored, enhancing the feasibility and attractiveness of CCS projects.
A cornerstone of the initiative is the inclusive and comprehensive development of infrastructure necessary for the effective transport and storage of CO2 across Poland. Conducting detailed feasibility assessments of various CO2 transport modalities—such as pipelines, railways, inland waterways, and export ports for international shipping facilities is paramount. The potential for repurposing existing pipelines or creating new routes along current corridors offers opportunities to optimize infrastructure usage.
These detailed feasibility studies, conducted in collaboration with the research and development (R&D) sector, assess technical viability while exploring a range of scenarios. This integration of diverse solutions forms a key component of the national CCS strategy, ensuring that infrastructure development is both strategic and forward-looking, thereby supporting Poland’s climate mitigation goals.
@@ 6.3 Public engagement and stakeholder participation
Active public engagement and stakeholder participation are indispensable for the successful deployment of CCS technologies (Fig. 5)). This engagement is not only about disseminating information but also about creating a continuous dialogue that addresses public concerns and builds trust. Drawing from current campaign experiences (discussed above) and stakeholder workshops, it is evident that education and awareness initiatives serve as fundamental tools in combating misinformation, which can otherwise hinder the acceptance and implementation of CCS projects.
A well-structured public campaign should leverage various communication channels to reach diverse demographic groups, emphasizing the importance of CCS in mitigating climate change and meeting emission reduction targets. This is particularly effective when initiated early, incorporating educational programs within school curriculums to instill an understanding of environmental challenges among younger generations. Such early engagement ensures a culture of awareness and responsibility, which is crucial for long-term support.
Public campaigns should also focus on explaining the technological processes, safety measures, and potential economic benefits of CCS. Demystifying the technology and showcasing successful examples, such as those in Norway and other regions, can boost public confidence. The use of relatable and accessible language in these campaigns will help bridge the gap between scientific complexities and public understanding.
Moreover, sustainable public support is fostered through transparency and the inclusion of public feedback in decision-making processes. Facilitating open dialogues with local communities and stakeholders not only resolves potential issues but also cultivates a sense of ownership and involvement among the populace. By addressing specific local concerns, such as environmental safety and economic impacts, these dialogues can remove barriers to acceptance.
Poland can adopt similar strategies based on Norway’s successful public communication practices, which have demonstrated efficacy in building stakeholder confidence. Norwegian experiences highlight the importance of harnessing trusted sources of information, such as scientific journals and academic channels, to disseminate accurate and reliable CCS data. Such strategic communication efforts ensure that public engagement is not a one-time initiative but an ongoing process that evolves with the project’s development.
@@ Acknowledgments
The authors acknowledge the support and funding received from the "Norwegian-Polish CCS Network: Acceleration of Climate Change Mitigation Technologies Deployment" project, funded through the EEA and Norway Grants under the Green Transition Collaborations Programme (FWD-Green-11). This research was facilitated by the collaborative efforts of the AGH University of Krakow, University of Oslo, and Norwegian Energy Partners (NORWEP).
CRediT authorship contribution
Mohammad Nooraiepour: Conceptualization, manuscript design, investigation, data curation and visualization, and writing the original draft. Pawel Gladysz: Writing – review & editing of final draft. Eirik Melaaen: Writing – review & editing of final draft.